# Through the Looking Glass: Mirror Schrödinger Bridges

### v1

@@ 4.3 Practical Algorithm
In this section, we describe an algorithm to solve the mirror Schrödinger bridge problem numerically, based on the AMP scheme that we introduced in section 4.1. We choose our reference path measure ${\mathbb{P}}^{0}\in{\mathbb{S}}$ to be associated to an Ornstein-Uhlenbeck process ${\mathbf{X}}_{t}$ given by an SDE of the form ${\rm{d}}{\mathbf{X}}_{t}=-\alpha{\mathbf{X}}_{t}{\rm{d}}t+\sigma{\rm{d}}{\mathbf{W}}_{t}$, for $\alpha>0$.
Algorithm 1 Mirror Schrödinger bridge
@@
Recall that our proposed AMP scheme alternates between direct $D_{\mathrm{KL}}$ projections on the set of path measures with a prescribed initial marginal distribution $\pi$ and reverse $D_{\mathrm{KL}}$ projections on the set of time-symmetric path measures. We now explain how each of these projections is computed in practice. Our algorithm then follows by iteratively applying this pair of projections and is summarized in Algorithm 1.
Direct $D_{\mathrm{KL}}$ projection. We can compute the $D_{\mathrm{KL}}$ projection onto the set of path measures with a prescribed initial marginal distribution $\pi$ following the trajectory-caching method developed and applied in (Vargas et al., 2021; De Bortoli et al., 2021). Let $\pi$ be a probability distribution on $\mathbb{R}^{n}$, and let ${\mathbb{P}}\in{\mathbb{D}}(\pi,\cdot)$ and ${\mathbb{P}}^{\dagger}\in{\mathbb{S}}$ be path measures corresponding to diffusion processes. Write $f_{t}^{{\mathbb{P}}}$ and $b_{t}^{{\mathbb{P}}}$ for the forward and backward drift functions corresponding to ${\mathbb{P}}$, and write $v_{t}^{{\mathbb{P}}^{\dagger}}=f_{t}^{{\mathbb{P}}^{\dagger}}=b_{t}^{{\mathbb{P}}^{\dagger}}$ for the drift of ${\mathbb{P}}^{\dagger}$. As a consequence of Girsanov’s theorem, we can write $D_{\mathrm{KL}}({\mathbb{P}}\;\|\;{\mathbb{P}}^{\dagger})$ explicitly in terms of $f_{t}^{{\mathbb{P}}}$ and $v_{t}^{{\mathbb{P}}^{\dagger}}$, or equivalently in terms of $b_{t}^{{\mathbb{P}}}$ and $v_{t}^{{\mathbb{P}}^{\dagger}}$; for references, see (Chen et al., 2016, section 3) as well as (Winkler et al., 2023, sections 2.2, 2.3). Indeed, for some constants $C_{1},\,C_{2}$, we have
$D_{\mathrm{KL}}({\mathbb{P}}\;\|\;{\mathbb{P}}^{\dagger})$ $=C_{1}+\frac{1}{2\sigma^{2}}\int_{0}^{1}\mathbb{E}_{{\mathbb{P}}}\left[\left(f_{t}^{{\mathbb{P}}}({\mathbf{X}}_{t})-v_{t}^{{\mathbb{P}}^{\dagger}}({\mathbf{X}}_{t})\right)\right]^{2}{\rm{d}}t$ (7)
$=C_{2}+\frac{1}{2\sigma^{2}}\int_{0}^{1}\mathbb{E}_{{\mathbb{P}}}\left[\left(b_{t}^{{\mathbb{P}}}({\mathbf{X}}_{t})-v_{t}^{{\mathbb{P}}^{\dagger}}({\mathbf{X}}_{t})\right)\right]^{2}{\rm{d}}t.$ (8)
In light of the identities (7) and (8), and because drift functions are much more amenable to modeling and estimation than path measures, it is convenient to recast the steps of our AMP scheme as iterative computations of drift functions associated to $D_{\mathrm{KL}}$ projections.
It follows immediately from (7) that the direct $D_{\mathrm{KL}}$ projection of ${\mathbb{P}}^{\dagger}$ onto the space ${\mathbb{D}}(\pi,\cdot)$ is given by the unique path measure ${\mathbb{P}}$ with initial marginal $\pi$ and forward drift $f_{t}^{{\mathbb{P}}}$ equal to the drift $v_{t}^{{\mathbb{P}}^{\dagger}}$ of ${\mathbb{P}}^{\dagger}$. In our AMP scheme, we employ this by taking ${\mathbb{P}}={\mathbb{P}}^{2k+1}\in{\mathbb{D}}(\pi,\cdot)$ to have drift equal to that of ${\mathbb{P}}^{\dagger}={\mathbb{P}}^{2k}$ for each $k\geq 0$. But as we will see in our analysis of the reverse $D_{\mathrm{KL}}$ projection, it does not suffice for us to know only the forward drift associated to our path measure iterates. We need to know the backward drift $b_{t}^{{\mathbb{P}}}$ too, but in practice, we do not have access to it. We use trajectory caching to estimate the backward drift $b_{t}^{{\mathbb{P}}}$. Trajectory caching is principled on the fact that $b_{t}^{{\mathbb{P}}}$ can be expressed in terms of the expected rate of change in ${\mathbf{X}}_{t}$ over time. Concretely, we have the following formula, which can be taken as a formal definition of the backward drift of a diffusion process:
@@
${\mathbf{X}}_{i+1}^{j}={\mathbf{X}}_{i}^{j}+f_{\bar{\gamma}_{i}}^{{\mathbb{P}}}({\mathbf{X}}_{i}^{j})\gamma_{i}+\sigma^{j}\sqrt{\gamma_{i}}{\mathbf{Z}}_{i}^{j},\quad\text{ where }\quad{\mathbf{Z}}_{i}^{j}\sim{\mathcal{N}}(0,{\bm{I}}).$ (10)
@@
The limiting quantity in (9) is then leveraged as the target of the loss function used to train a neural network $v_{t}^{\theta}$, which approximates the backward drift $b_{t}^{\mathbb{P}}$ for a specified range of $\sigma$ values $[\sigma_{\min},\sigma_{\max}]$. Specifically, we define the following loss function in terms of the optimization parameter $\theta$:
@@
$\ell(\theta)=\frac{1}{N}\sum_{i=1}^{M-1}\sum_{j=0}^{N-1}\left|\left|v_{\bar{\gamma}_{i+1}}^{\theta}({\mathbf{X}}_{i+1}^{j})-\frac{{\mathbf{X}}_{i}^{j}-{\mathbf{X}}_{i+1}^{j}}{\gamma_{i+1}}-\left(f_{\bar{\gamma}_{i}}^{{\mathbb{P}}^{\dagger}}({\mathbf{X}}_{i+1}^{j})-f_{\bar{\gamma}_{i}}^{{\mathbb{P}}^{\dagger}}({\mathbf{X}}_{i}^{j})\right)\right|\right|^{2}$ (11)
@@
Observe that the first two terms in the loss constitute the difference between the drift and the infinitesimal rate of change of the process ${\mathbf{X}}_{t}$, i.e., the discretization of the difference between the left- and right-hand sides of (9). The network parameters $\theta$ are then learned via gradient descent with respect to the loss function $\ell(\theta)$. The resulting function $v_{t}^{\theta}$, where $\theta$ minimizes the loss $\ell(\theta)$, approximates the desired backward drift, as is suggested by De Bortoli et al. (2021, Proposition 3).
Reverse $D_{\mathrm{KL}}$ projection. We now describe how to compute the reverse $D_{\mathrm{KL}}$ projection onto the set ${\mathbb{S}}$ of time-symmetric path measures. We are interested in computing the associated time-symmetric drift, rather than the path measure itself. To this end, let $\pi$ be a probability distribution on $\mathbb{R}^{n}$, and let ${\mathbb{P}}\in{\mathbb{D}}(\pi,\cdot)$ and ${\mathbb{P}}^{\dagger}\in{\mathbb{S}}$ be path measures corresponding to diffusion processes. Suppose we seek to minimize $D_{\mathrm{KL}}({\mathbb{P}}\;\|\;{\mathbb{P}}^{\dagger})$ over all ${\mathbb{P}}^{\dagger}\in{\mathbb{S}}$. Using (7) and (8), we can write $D_{\mathrm{KL}}({\mathbb{P}}\;\|\;{\mathbb{P}}^{\dagger})$ explicitly in terms of the forward and backward drift functions of the SDE corresponding to the path measures ${\mathbb{P}}$ and ${\mathbb{P}}^{\dagger}$. A key benefit of considering the reverse $D_{\mathrm{KL}}$ projection is that the expectation values in (7) and (8) are taken with respect to the fixed path measure ${\mathbb{P}}$, and not with respect to the varying path measure ${\mathbb{P}}^{\dagger}$. This allows us to apply calculus of variations to compute a closed-form expression for the drift of the minimizer of $D_{\mathrm{KL}}({\mathbb{P}}\;\|\;{\mathbb{P}}^{\dagger})$ over ${\mathbb{P}}^{\dagger}\in{\mathbb{S}}$. First, observe that we can combine (7) and (8) to rewrite $D_{\mathrm{KL}}({\mathbb{P}}\;\|\;{\mathbb{P}}^{\dagger})$ in a time-symmetric formulation as follows:
$D_{\mathrm{KL}}({\mathbb{P}}\;\|\;{\mathbb{P}}^{\dagger})$ $=C+\frac{1}{4\sigma^{2}}\int_{0}^{1}{\mathbb{E}_{{\mathbb{P}}}}\left[\left(f_{t}^{{\mathbb{P}}}({\mathbf{X}}_{t})-v_{t}^{{\mathbb{P}}^{\dagger}}({\mathbf{X}}_{t})\right)^{2}+\left(b_{t}^{{\mathbb{P}}}({\mathbf{X}}_{t})-v_{t}^{{\mathbb{P}}^{\dagger}}({\mathbf{X}}_{t})\right)^{2}\right]{\rm{d}}t,$
where $C$ is a constant. Note that the sum of squares inside the expectation on the right-hand side above is always nonnegative. Consequently, to minimize $D_{\mathrm{KL}}({\mathbb{P}}\;\|\>{\mathbb{P}}^{\dagger})$, it suffices to choose ${v_{t}^{{\mathbb{P}}}}^{\dagger}$ so that it minimizes this sum of squares pointwise everywhere. Taking the first variation of this sum of squares with respect to $v_{t}^{{\mathbb{P}}^{\dagger}}$, setting the resulting expression equal to zero, and solving for the optimal $v_{t}^{{\mathbb{P}}^{\dagger}}$, we find that
That is, the choice of ${\mathbb{P}}^{\dagger}\in{\mathbb{S}}$ minimizing $D_{\mathrm{KL}}({\mathbb{P}}\;\|\;{\mathbb{P}}^{\dagger})$ has drift function given by the average of the forward and backward drifts of ${\mathbb{P}}$.
In our AMP scheme, we employ (12) by taking the forward and backward drifts corresponding to ${\mathbb{P}}={\mathbb{P}}^{2k+1}$ and averaging them to obtain the drift of the symmetric path measure ${\mathbb{P}}^{\dagger}={\mathbb{P}}^{2k+2}$. Practically speaking, if the drift of ${\mathbb{P}}^{2k}$ is a parametrized by a neural network $v_{t}^{\theta^{k}}$ for each $k$, we take the drift of ${\mathbb{P}}^{2k+1}$ to be the average of the outputs of the neural networks $v_{t}^{\theta^{k-1}}$ and $v_{t}^{\theta^{k}}$. In Algorithm 1, we denote the limiting drift as $v_{t}^{\theta^{\star}}$.
4.4 Sampling with In-distribution Variation
In this section, we provide a short intuitive explanation of how our method allows for resampling with prescribed proximity to an input sample. Given such a sample $x_{0}\sim\pi$, we solve the SDE corresponding to the Schrödinger bridge to push $x_{0}$ forward in time, arriving at a final sample $x_{1}\in\pi$. We want $x_{1}$ to be a variation of $x_{0}$, where the proximity of $x_{1}$ to $x_{0}$ correlates with the size of the noise coefficient $\sigma$. Justifying this mathematically requires understanding how the conditional distribution ${\mathbf{X}}_{1}\mid{\mathbf{X}}_{0}=x_{0}$, specifically its mean and variance, depend on $\sigma$. While these quantities do not in general have closed form expressions, it is possible to compute them exactly in the case where $\pi={\mathcal{N}}(0,{\mathbf{I}})$ is a $1$-dimensional Gaussian.
In this case, let ${\mathbf{X}}_{t}$ denote the diffusion process associated to the Schrödinger bridge, where the reference path measure corresponds to an Ornstein-Uhlenbeck reference process with drift coefficient $-\alpha$. In Proposition 4 (see Appendix B for the statement and proof) we determine the joint distribution of ${\mathbf{X}}_{0}$ and ${\mathbf{X}}_{1}$ in terms of a quantity $\beta$, which is a function of $\alpha$ and $\sigma$ that grows approximately as $1+c(\alpha)\times\sigma^{2}$ for some function $c$. Let $p(x,y)$ denote the probability density function of the joint distribution of ${\mathbf{X}}_{0}$ and ${\mathbf{X}}_{1}$, and recall that $p(x,y)$ is the product of the conditional PDF of ${\mathbf{X}}_{1}\mid{\mathbf{X}}_{0}$ with the PDF of ${\mathbf{X}}_{0}$. Using this fact in conjunction with Proposition 4, the PDF of ${\mathbf{X}}_{1}\mid{\mathbf{X}}_{0}=x_{0}$ is
$p_{{\mathbf{X}}_{1}\mid{\mathbf{X}}_{0}=x_{0}}(y)=\nicefrac{{p(x_{0},y)}}{{p_{{\mathbf{X}}_{0}}(x_{0})}}=e^{-\frac{1}{2(1-\beta^{2})}(x_{0}^{2}-2\beta x_{0}y+y^{2})+\frac{x_{0}^{2}}{2}}.$
From the right-hand side, we see that ${\mathbf{X}}_{1}\mid{\mathbf{X}}_{0}=x_{0}$ is Gaussian with mean and variance given by
$\mathbb{E}[{\mathbf{X}}_{1}\mid{\mathbf{X}}_{0}=x_{0}]=x_{0}\left(\nicefrac{{\beta}}{{1-\beta^{2}}}\right),\;\mathbb{E}\big{[}({\mathbf{X}}_{1}-\mathbb{E}[{\mathbf{X}}_{1}\mid{\mathbf{X}}_{0}=x_{0}])^{2}\mid{\mathbf{X}}_{0}=x_{0}\big{]}=1-\beta^{2}.$
Thus, changing the noise value $\sigma$ alters both the mean and variance of samples pushed forward via the Schrödinger bridge. Indeed, in the case of the mean, it grows inversely proportional to $\sigma^{2}$. Consequently, if $\sigma<1$, then we should expect the Schrödinger bridge to push samples away from the distribution mean, whereas if $\sigma>1$, then the opposite occurs, and samples experience mean reversion. As for the variance, note that $1-\beta^{2}$ grows at least as fast as $\sigma^{2}$, so we should expect the Schrödinger bridge to produce samples with spread that increases as $\sigma$ increases. We expect that similar effects occur even when the marginal distribution $\pi$ is not Gaussian: i.e., the value of $\sigma$ should be directly related to the proximity of generated samples in an analogous way.
@@ 5 Experiments
We demonstrate the flexibility of our method on a number of conditional sampling tasks. We first show numerical convergence against the solution of the mirror Schrödinger bridge in a case where an analytical solution is available. Next, we consider resampling from $2$-dimensional datasets and demonstrate control over the in-distribution variation of new data points, which is an added feature of our method. Lastly, we provide examples of image resampling, illustrating how our method can be used to produce image variations with control over the proximity to the original.
5.1 Gaussian Transport
@@
We start by comparing our method with two alternative algorithms, DSB (De Bortoli et al., 2021) and DSBM (Shi et al., 2023), when applied to the mirror Schrödinger bridge case on Gaussians of varying dimension. Figure 1 shows that, in the case of dimension $d=50$, as the number of outer iterations increases, the empirical convergence of our method performs on par with both DSB and DSBM with the added benefit that each outer iteration with our algorithm requires half the training iterations. Recall that our method trains a single neural network to model a time-symmetrized drift function $v_{t}^{\theta}$ rather than a neural network for each of the forward and backward drift functions. More details on the derivation of the analytical solution for this experiment, as well as information on parameters, can be found in Appendix B. Additional results for dimensions $d=5,20$ can be found in Figure 6.
5.2 2D Datasets
@@
To illustrate the behavior of our method, we use our algorithm to resample from 2-dimensional distributions. Unlike the mirror Schrödinger bridge with Gaussians, an analytical solution for mirror bridge with these more general distributions is not known. We consider learning the drift function $v_{t}^{\theta}$ associated with the mirror Schrödinger bridge that flows samples from $p_{\rm{data}}$ to itself. The goal is to obtain new samples that are in the distribution $p_{\rm{data}}$ but exhibit some level of variation, i.e., in-distribution variation, correlated to the noise coefficient $\sigma$ in the diffusion process. Note that computing mirror Schrödinger bridges with a range of noise values by training one neural network is not possible using existing alternative methods.
Figure 2: Samples obtained using our method with varying values of $\sigma$. Samples are colored based on initial position.
@@
In the columns of Figure 2, we show the result of flowing samples via the mirror Schrödinger bridge with varying values of noise. We observe that the in-distribution variation of data points is controlled by the choice of $\sigma$ value, which can indeed be detected by the mixing of colors, or lack of thereof, in each terminal distribution shown. For instance, in the bottom row, we find mixing from samples between the inner and outer circles with the largest value of $\sigma$, compared with no mixing of samples between circles with the smallest value of sigma.
(a) Initial samples
(b) $\sigma=1$
(c) $\sigma=2$
Figure 3: Samples produced by the mirror Schrödinger bridges for the empirical distribution of handwritten digits, using varying levels of noise.
5.3 Image Resampling
@@
We also train our algorithm on each of the MNIST, CelebA, and Flower102102102102 datasets. Details on training parameters and architecture for all experiments using images can be found in Appendix C. Our results show that mirror Schrödinger bridges can be used to produce new samples from an image dataset with control over the proximity to the initial sample. In Figure 3, we resample from MNIST using varying levels of noise. We find that pushforward images obtained with a lower fixed value of noise (3(b)) are visually closer to the initial images (3(a)) obtained with a higher fixed value of noise (3(c)).
@@
Figure 4 demonstrates the same control over the in-distribution variation of pushforward samples using the RGB dataset CelebA. In each column, we exhibit a different sample from the dataset and, from top to bottom, we show the pushforward samples obtained with varying levels of noise. Across a fixed column, all images resemble the initial sample to varying degrees. In particular, note that the images on the bottom are the least similar to the initial sample, whereas the images on the second row are the most proximal. Images across rows were obtained without retraining the network. This can be done as long as the values chosen for the noise $\sigma$ are within the range of values on which the network was trained. Figure 8 includes more results using the CelebA dataset, and Figure 7 shows the nearest neighbors in the dataset to the generated images. In the latter Figure, as desired, the nearest neighbor of the generated sample is the initial sample itself, and the generated sample is distinct from all of its nearest neighbors, showing that our model does not simply regurgitate nearest neighbors of the initial sample as proximal outputs.
@@
Figure 4: The control over in-distribution variance effect of $\sigma$ for a variety of initial samples (first row) from the empirical distribution of images in CelebA.
Figure 5 highlights how mirror Schrödinger bridges can be used as a flexible and well-principled tool to perform small edits to RGB images while guaranteeing the result to be in-distribution. This task can be performed by choosing an appropriately small value for $\sigma$.

### v2

@@ 4.3 Practical Algorithm
Algorithm 1 Mirror bridge
@@
We now describe an algorithm to solve the mirror bridge problem numerically, based on our AMP scheme from §4.2. We choose our reference path measure ${\mathbb{P}}^{0}\in{\mathbb{S}}$ to be associated to an Ornstein-Uhlenbeck process ${\mathbf{X}}_{t}$ given by the ${\rm{d}}{\mathbf{X}}_{t}=-\alpha{\mathbf{X}}_{t}{\rm{d}}t+\sigma{\rm{d}}{\mathbf{W}}_{t}$, for $\alpha>0$.
Recall that our AMP scheme alternates between direct $D_{\mathrm{KL}}$ projections on the set of path measures with initial marginal $\pi$ and reverse $D_{\mathrm{KL}}$ projections on the set of time-symmetric path measures. We now explain how each of these projections is computed in practice. Our algorithm then follows by iteratively applying this pair of projections and is summarized in Algorithm 1.
As is evident from our analysis of the reverse $D_{\mathrm{KL}}$ projection, it does not suffice for us to know only the forward drift associated to our path measure iterates. We need to know the backward drift $b_{t}^{{\mathbb{P}}}$ too, but in practice, we do not have access to it. We use trajectory caching to estimate the backward drift $b_{t}^{{\mathbb{P}}}$. Trajectory caching is principled on the fact that $b_{t}^{{\mathbb{P}}}$ can be expressed in terms of the expected rate of change in ${\mathbf{X}}_{t}$ over time. Concretely, we have the following formula, which can be taken as a formal definition of the backward drift of a diffusion process:
@@
${\mathbf{X}}_{i+1}^{j}={\mathbf{X}}_{i}^{j}+f_{\bar{\gamma}_{i}}^{{\mathbb{P}}}({\mathbf{X}}_{i}^{j})\gamma_{i}+\sigma^{j}\sqrt{\gamma_{i}}{\mathbf{Z}}_{i}^{j},$ (13)
@@
where ${\mathbf{Z}}_{i}^{j}\sim{\mathcal{N}}(0,{\bm{I}})$. The limiting quantity in (12) is then leveraged as the target of the loss function used to train a neural network $v_{t}^{\theta}$, which approximates the backward drift $b_{t}^{\mathbb{P}}$ for a specified range of $\sigma$ values $[\sigma_{\min},\sigma_{\max}]$. Specifically, we define the loss $\ell$ in terms of the optimization parameter $\theta$ by
@@
$\ell=\frac{1}{N}\sum_{i=1}^{M-1}\sum_{j=0}^{N-1}\bigg{|}\bigg{|}v_{\bar{\gamma}_{i+1}}^{\theta}({\mathbf{X}}_{i+1}^{j})-\frac{{\mathbf{X}}_{i}^{j}-{\mathbf{X}}_{i+1}^{j}}{\gamma_{i+1}}-\left(f_{\bar{\gamma}_{i}}^{{\mathbb{P}}^{\dagger}}({\mathbf{X}}_{i+1}^{j})-f_{\bar{\gamma}_{i}}^{{\mathbb{P}}^{\dagger}}({\mathbf{X}}_{i}^{j})\right)\bigg{|}\bigg{|}^{2}$ (14)
@@
Observe that the first two terms in the loss constitute the difference between the drift and the infinitesimal rate of change of the process ${\mathbf{X}}_{t}$, i.e., the discretization of the difference between the left- and right-hand sides of (12). The network parameters $\theta$ are then learned via gradient descent with respect to the loss function $\ell(\theta)$. The resulting function $v_{t}^{\theta}$, where $\theta$ minimizes the loss $\ell(\theta)$, approximates the desired backward drift, as is suggested by De Bortoli et al. (2021, Proposition 3). In our AMP scheme, we employ (11) by averaging the forward and backward drifts corresponding to ${\mathbb{P}}={\mathbb{P}}^{2k+1}$ to obtain the drift of the symmetric path measure ${\mathbb{P}}^{\dagger}={\mathbb{P}}^{2k+2}$. Practically speaking, if the drift of ${\mathbb{P}}^{2k}$ is a parametrized by a neural network $v_{t}^{\theta^{k}}$ for each $k$, we take the drift of ${\mathbb{P}}^{2k+1}$ to be the average output of the networks $v_{t}^{\theta^{k-1}}$ and $v_{t}^{\theta^{k}}$. In Algorithm 1, we denote the limiting drift as $v_{t}^{\theta^{\star}}$.
Sampling with in-distribution variation. We provide a short intuitive explanation of why our method allows for resampling with prescribed proximity to an input sample in Appendix C.
@@
Figure 3: Samples (color based on initial position) obtained using our method with various $\sigma$ values.
@@ 5 Experiments
We demonstrate the utility of our method on a number of conditional resampling tasks, illustrating how it can be used to produce sample variations with control over the proximity to the initial sample.
@@
Gaussian transport. We start by comparing our method with two alternative algorithms, DSB (De Bortoli et al., 2021) and DSBM (Shi et al., 2023), when applied to the mirror Schrödinger bridge of multi-dimensional Gaussians. Fig. 2 shows that, in the case of dimension $d=50$, as the number of outer iterations increases, the empirical convergence of our method performs on par with both DSB and DSBM with the added benefit that each outer iteration with our algorithm requires half the training iterations. Recall that our method trains a single neural network to model a time-symmetrized drift function $v_{t}^{\theta}$ rather than a neural network for each of the forward and backward drift functions. More details on the derivation of the analytical solution, as well as information on parameters, can be found in Appendix D. Additional results for dimensions $d=5,20$ can be found in Fig. 6.
@@
2D datasets. To illustrate the behavior of our method, we use our algorithm to resample from 2-dimensional distributions Pedregosa et al. (2011). Unlike the mirror Schrödinger bridge with Gaussians, an analytical solution with these more general distributions is not known. We consider learning the drift function $v_{t}^{\theta}$ associated with the mirror bridge that flows samples from $p_{\rm{data}}$ to itself. The goal is to obtain new samples that are in the distribution $p_{\rm{data}}$ but exhibit some level of variation, i.e., in-distribution variation, correlated to the noise coefficient $\sigma$ in the diffusion process.
@@
In Fig. 3, we show the result of flowing samples via the mirror bridge with varying values of noise. We observe that the in-distribution variation of data points is controlled by the choice of $\sigma$ value, which can indeed be detected by the mixing of colors, or lack of thereof, in each terminal distribution shown. For instance, on the right, we find mixing from samples between the inner and outer circles with the largest $\sigma$ value, compared with no mixing between circles with the smallest $\sigma$ value.
@@
Image resampling. We train our algorithm on the MNIST LeCun et al. (1998) and CelebA Liu et al. (2015) datasets. Training parameters and architecture for all experiments can be found in Appendix E. Our results show that mirror bridges can be used to produce new samples from an image dataset with control over the proximity to the initial sample. In Fig. 4, we resample from MNIST using varying levels of noise (see also Fig. 11). Pushforward images obtained with a lower $\sigma$ value (middle) are visually closer to the initial images (top) than the ones obtained with a higher $\sigma$ value (bottom).
Figure 4: Samples produced by our mirror bridges for MNIST, using $\sigma=1,2$.
@@
Fig. 5 shows the same control over in-distribution variation of pushforward samples using the RGB dataset CelebA. In each column, we exhibit a different sample from the dataset and, in each row, we show the corresponding pushforward obtained for different noise values. These results can be obtained without retraining the neural network. The typical metric to assess resampling quality for the image generation case is the Fréchet inception distance (FID) score, which we have plotted against training iterations. We observe FID scores decreasing with training iterations. Fig. 12 includes more results using the CelebA dataset, and Fig. 7 shows the nearest neighbors in the dataset to the generated images. In the latter, the nearest neighbor of the generated sample is the initial sample itself, and the generated sample is distinct from all of its nearest neighbors, showing that our model does not simply regurgitate nearest neighbors of the initial sample as proximal outputs. Additional experiments, including a comparison to alternative methods for image resampling, further evaluation metrics, results on path regularity and control over proximity, are presented in Appendix F.
@@
Figure 5: The control over in-distribution variance effect of $\sigma$ for a variety of initial samples (first row) from the empirical distribution of images in the CelebA dataset.